OpenStudy (anonymous):

what is the simplified form of... (25x^8/24y^5) divided by (4x^4/8y^2)

sam (.sam.):

((25x^(8))/(24y^(5)))/((4x^(4))/(8y^(2))) (8y^(2))/(4x^(4))*(25x^(8))/(24y^(5)) (2y^(2))/(x^(4))*(25x^(8))/(24y^(5)) 1*(25x^(4))/(12y^(3)) 25x^(4)/12y^(3)

OpenStudy (anonymous):

it has to be simplified more that not one of my choices

OpenStudy (callisto):

\[\frac{\frac{25x^8}{24y^5}}{\frac{4x^4}{8y^2}}\]\[=\frac{25x^8}{24y^5}\times \frac{8y^2}{4x^4}\]\[=\frac{25x^8}{24y^5}\times \frac{2y^2}{x^4}\]\[=\frac{25x^4}{12y^3}\]